OpenStudy (anonymous):

Which of the following is an arithmetic sequence? 2, 4, 8, 16, ... 12, 4, 4/3, 16/3 ... 1/2, -1/2, -3/2, -5/2 ... 1/2, -1/2, 1/2, -1/2 ...

OpenStudy (mathstudent55):

Do you know the difference between an arithmetic sequence and a geometric sequence?

OpenStudy (anonymous):

No, I haven't learned about geometric sequence yet.

OpenStudy (mathstudent55):

In an arithmetic sequence each term minus the previous term is always the same difference. That is called the common difference.

OpenStudy (anonymous):

arithmetic sequence is A sequence made by adding the same value each time.

OpenStudy (mathstudent55):

Ok. Then we will only discuss the arithmetic sequence. Start with the second term, and subtract the first term from it. You get a difference. Then subtract the second term from the third term. Subtract the third term from the fourth term. If the difference is always the same, that is the common difference and it is an arithmetic sequence.

OpenStudy (anonymous):

Okay so how should I determine he answer?

OpenStudy (mathstudent55):

Let's start with the first choice. You need to do this to all the choices, but we'll use the first choice to do together.

OpenStudy (anonymous):

Okay.

OpenStudy (mathstudent55):

Subtract the first term from the second term: what is 4 - 2 = ?

OpenStudy (anonymous):

2

OpenStudy (mathstudent55):

Good. Now subtract the second term form the third term. What is 8 - 4 = ?

OpenStudy (anonymous):

4

OpenStudy (mathstudent55):

Good. The two differences are different. You don't have a common difference, therefore the first choice is NOT an arithmetic sequence.

OpenStudy (anonymous):

Okay.

OpenStudy (mathstudent55):

Remember, in an arithmetic, the difference between any term and the previous term is always the same.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Thank you for your help!... I needed that.

OpenStudy (mathstudent55):

Now try choice B. Do the same thing. What is 4 - 12 = ?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

The fractions are what gets me.

OpenStudy (mathstudent55):

Good. Then do 4/3 - 4 = ? If these two differences are not equal, it is not an arithmetic sequence.

OpenStudy (mathstudent55):

Then do the same for the other 2 choices.

OpenStudy (anonymous):

okay. Thank you!